# DreamMapping: High-Fidelity Text-to-3D Generation via Variational Distribution Mapping

### v1

@@ Abstract
Score Distillation Sampling (SDS) has emerged as a prevalent technique for text-to-3D generation, enabling 3D content creation by distilling view-dependent information from text-to-2D guidance. However, they frequently exhibit shortcomings such as over-saturated color and excess smoothness. In this paper, we conduct a thorough analysis of SDS and refine its formulation, finding that the core design is to model the distribution of rendered images. Following this insight, we introduce a novel strategy called Variational Distribution Mapping (VDM), which expedites the distribution modeling process by regarding the rendered images as instances of degradation from diffusion-based generation. This special design enables the efficient training of variational distribution by skipping the calculations of the Jacobians in the diffusion U-Net. We also introduce timestep-dependent Distribution Coefficient Annealing (DCA) to further improve distilling precision. Leveraging VDM and DCA, we use Gaussian Splatting as the 3D representation and build a text-to-3D generation framework. Extensive experiments and evaluations demonstrate the capability of VDM and DCA to generate high-fidelity and realistic assets with optimization efficiency. {CCSXML} <ccs2012> <concept> <concept_id>10010147.10010178.10010224.10010226.10010239</concept_id> <concept_desc>Computing methodologies 3D imaging</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010382.10010383</concept_id> <concept_desc>Computing methodologies Image processing</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
@@ 1 Introduction
There has been an increasing need for digital 3D assets in video games [ZW12], mixed reality storytelling [CT14, HWW21], and digital fabrications [Ger12]. However, this widespread growth highlights the need for more efficient 3D content creation, as crafting 3D assets in graphics software is often time-consuming and labor-intensive. Thanks to the rapid development of diffusion models [HJA20, DN21], recent text-to-3D techniques offer a more accessible and efficient solution for controllable 3D content creation.
Based on the pre-trained text-to-image models [RBL21, SME20], Score Distillation Sampling (SDS) [PJBM22, WDL23] provides pixel-level guidance of 3D assets optimization and has become quite effective and popular for text-to-3D generation. Follow-up works further refine the supervision loss and optimization procedures to produce more realistic and semantic-aligned 3D models [WLW23, CCH23, CCJJ23, YGL23, KPCOL23, WZY24, LSS24, AKS24, WZSZ23, LYL23, TWWZ23, YCC23].
In this paper, we review SDS-based 3D generation and its variations, leading to a unifying goal across previous work on refining SDS formulation: establishing a variational distribution for rendered images. Our review and analysis demonstrate that prior efforts have predominantly adopted one of two strategies: 1) utilizing unmodified diffusion models to approximate the variational distribution [YGL23, KPCOL23, LYL23, WZY24] with improved loss designs, and 2) constructing the variational distribution directly through fine-tuning diffusion models [WLW23, LSS24, WZSZ23, YCC23] or pre-training a bespoke model [AKS24]. However, the former approach struggles to represent the distribution accurately, since the rendered images during the optimization process are out-of-domain (OOD) cases as to diffusion models [KPCOL23]. Meanwhile, the latter choice demands substantial resources and may exhibit instability during optimization.
To address these issues, we propose variational distribution mapping (VDM), a novel method treating rendered images as a degraded form of images generated by the diffusion model. VDM efficiently formulates the variational distribution of rendered images by modeling and optimizing the degradation process (i.e., lightweight trainable neural network), which enables us to establish a distribution mapping between the generated and the rendered image distributions (Figure 1). Compared with previous methods, VDM has two notable advantages: 1) By introducing the trainable degradation process, it eliminates the need for the complex Jacobian matrix calculations in the UNet of diffusion model, unlike previous variational distribution modeling approaches [WZSZ23, YCC23]; 2) VDM dynamically models the variational distribution of rendered images, surpassing methods that rely on an unmodified diffusion model with extra batch processing [KPCOL23, YGL23] for distribution estimation, thereby enhancing generation quality.
@@ 2.1 Text-to-3D Generation
With the rapid advancement of diffusion models and 3D representations, significant progress has been made in text-based generation. DreamFields [JMB22] introduced the optimization of Neural Radiance Fields (NeRF) using guidance from pre-trained CLIP models, but the quality was limited due to the CLIP loss’s insufficient semantic guidance. Building on this, DreamFusion [PJBM22] and SJC [WDL23] developed SDS, a novel loss function based on probability density distillation that provides pixel-level guidance by seeking specific modes in a text-guide diffusion model, enhancing the quality and efficiency of optimization-based 3D generation.
The advent of SDS marked a turning point that inspired a series of subsequent works in 3D generation [WLW23, YGL23, KPCOL23, LSS24, LYL23, YCC23, WZSZ23]. These studies have refined the generation of 3D assets from various perspectives. Some utilizes advanced differentiable 3D representations to enhance outcomes [LGT23, CCJJ23], while others [WZY24]focus on mitigating the “Janus” problem [HAK23, AKS24]. Zero-1-to-3 [LWH23] and MVDream [SWY23] fine-tuned the diffusion model with multi-view image datasets to improve 3D consistency. Perp-Neg [ASZ23] alleviates the Janus problem with view-dependent negative prompts, and ESD [WXF23] introduces a view-conditioned loss to improve multi-view generation.
A central emphasis in recent research lies in refining the original SDS methodology, aiming to secure more precise guidance from diffusion models. ProlificDreamer [WLW23] introduces a Variational Score Distillation Sampling technique to alleviate the mode collapse issue inherent to SDS. To further mitigate over-saturation and over-smoothness problems, TextMesh [TMT24] enforces textured optimizations with high SDS gradients only and integrates multi-view consistent diffusion. Make-it-3D [TWZ23] proposes two-stage optimizations to enhance textured appearance; Fantasia3D [CCJJ23] dynamically adjusts the time-dependent weighting function in SDS computations. Meanwhile, CSD [YGL23] and NFSD [KPCOL23] demonstrate the pivotal role of the conditional term in SDS for 3D generation. They employ additional negative prompts to refine the optimization process. Their explorations highly inspired our work, and we propose further improving SDS with the trainable construction of variational distribution. We will continue more in-depth discussions in Sec. 3.1.
@@ 2.2 Differentiable 3D Representations
The differentiable 3D representations are pivotal to text-to-3D generation. A 3D model parameterized by $\theta$ is rendered into an image from camera viewpoint $c$ using the differentiable rendering function $g(\theta,c)$. This allows for optimizing the 3D models through backpropagation to align with pixel-level guidance derived from diffusion models. Numerous differentiable 3D representations have been utilized in prior text-to-3D works [PJBM22, LGT23, MRP22, GAG23, TRZ23], including NeRF [MST20] which is frequently employed in generation tasks. However, NeRF’s volume rendering is computationally intensive, posing challenges for rendering high-resolution images efficiently enough to leverage the diffusion model’s guidance.
To mitigate this constraint, DMTet [SGY21] has been introduced to integrate both explicit and implicit representations, which has gained increasing utilizations [CCJJ23, LWW23]. Additionally, there is a growing interest in applying purely explicit representations to facilitate smooth shape manipulations within graphics software. This involves a gradient-based mesh deformation approach to create 3D models [GAG23, WMC24]. Notably, 3D Gaussian Splatting [KKLD23] has emerged as an efficient and high-quality explicit representation for reconstruction tasks and has been incorporated into several text-to-3D generation works [TRZ23, YFW23, LYL23]. We investigate Gaussian Splatting as the 3D representation for our generation framework.
Preliminaries
@@ 2.3 Diffusion Models
The Diffusion Model [HJA20, SME20, SSDK20] is a likelihood-based generative model designed to approximate the data distribution $p_{data}$ starting from gaussian noise. Given a sample $\mathbf{x}$ from $p_{data}$, the model undergoes a forward diffusion process over a series of timesteps $t\in[0,t_{max}]$, whereby noise is incrementally added to transform $\mathbf{x}$ into Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$. The diffusion at each step is described by:
@@
The training goal is to optimize $\mu_{\phi}(\mathbf{x}_{t})$ and $\Sigma_{\phi}(\mathbf{x}_{t})$ to maximize the log likelihood’s variational lower bound. In practice, the learning target is re-parameterized to the added noise $\epsilon\sim\mathcal{N}(0,I)$ which is used to produce $\mathbf{x}_{t}$ from $\mathbf{x}=\mathbf{x}_{0}$, and the notation of UNet in diffusion models is $\epsilon_{\phi}(\mathbf{x}_{t},t)$. The learning process of diffusion models can be interpreted as predicting the noise $\epsilon$ that corrupts the data $\mathbf{x}$. Furthermore, recent studies [SSDK20, SE19] have shown that $\epsilon_{\phi}(\mathbf{x}_{t},t)$ corresponds to the score function $\nabla_{\mathbf{x}_{t}}\log p_{\phi}(\mathbf{x}_{t})$, indicating the model’s ability to guide $\mathbf{x}_{t}$ towards regions of higher density within $p_{\phi}(\mathbf{x_{t}})$, which can be approximated as:
@@
Classifier Free Guidance: Compared with unconditional image generation, text-guided image generation [HS21, SCS22, RBL21] has a higher demand. Conditioned on the text prompt $y$, diffusion models accept it as another input for the diffusion process, denoted as $\epsilon_{\phi}(\mathbf{x}_{t},t,y)$, with related score function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|y)$. Classifier Free Guidance (CFG) [HS21] is utilized as an implicit classifier to get the textual guidance for image generation. It has a changeable hyperparameter named CFG scale, hereafter denoted as $s$, and the original prediction is changed as:
@@ 2.4 SDS with Differentiable Rendering
As demonstrated in Sec. 2.1, SDS [PJBM22, WDL23] is an optimization-driven 3D generative technique that leverages pre-trained 2D diffusion models. Let $\mathbf{x}_{0}=g(\theta,c)$ represent the rendered image derived from differentiable 3D representations, parameterized by $\theta$ and camera pose $c$. The distribution of noisy rendered images is then formulated as follows:
@@
SDS adopts text-conditioned noisy real image distribution $p_{\phi}(\mathbf{x}_{t}|y)$ represented by pre-trained Stable Diffusion [RBL21] and optimizes the parameter $\theta$ by minimizing the following KL divergence for all timestep $t$:
@@
where $\epsilon\sim\mathcal{N}(0,I)$ and the UNet Jacobian term $\frac{\partial\epsilon_{\phi}(\mathbf{x}_{t},t,y)}{\partial\mathbf{x}_{t}}$ is ignored based on the analysis in DreamFusion [PJBM22].
@@
SDS [PJBM22] $\epsilon_{\phi}(\mathbf{x}_{t},t,y)-\epsilon$ Distillation baseline
@@
NFSD [KPCOL23] $\epsilon_{\phi}(\mathbf{x}_{t},t,y)-\lambda_{t}\epsilon_{\phi}(\mathbf{x}_{t},t,y_{neg})$
@@ 3.1 Systematic Review of Advancements in SDS
which delineates two components: the unconditional and the conditional terms. Numerous studies have indicated that the conditional term pivotal in ensuring the rendered image aligns with textual semantics [WLW23, YGL23, KPCOL23, TWWZ23]. This term is further guided by the score function $\nabla_{\mathbf{x}_{t}}\log p_{\phi}(y|\mathbf{x}_{t})$, as indicated in Eq. (5). Consequently, subsequent research mostly focuses on refining the unconditional term to improve SDS.
@@
Numerous studies have proposed various strategies to address the challenge of mode-seeking in 3D generation. Our review of recent advancements in refining the standard SDS approach reveals that these efforts generally focus on creating a variational distribution for the rendered images, which shifts the mode-seeking operation to minimizing the KL divergence between two distinct distributions. As illustrated in Table 1, VSD [WLW23] uses text-conditioned particles (or rendered images when the number of particles equals one) to build a variational distribution of 3D representations. VSD employs a LoRA [HWAZ21] model to swiftly optimize the score of rendered images, termed as $\epsilon_{\text{lora}}(\mathbf{x}_{t},t,y)$. DreamFlow [LSS24] follows VSD with a similar design, guided by theories of Schrodinger Bridge [LVH23]. ASD [WZSZ23] enhances the optimization methods of VSD within the framework of GAN [GPAM14], while LODS [YCC23] diversifies the modeling methods through learnable text embeddings denoted as $y_{\psi}$. LMC-SDS [AKS24] pre-trains a neural network to learn the score manifold corrective transitioning rendered images to generated images. The aforementioned studies collectively endeavor to determine the variational distribution of rendered images directly. Despite employing parameter-efficient modeling techniques, i.e., LoRA, the optimization of variational distribution requires additional computation of the UNet Jacobian for Stable Diffusion, which is time-consuming and prone to instability. For instance, given a text prompt, VSD, ASD, and LODs require approximately 8 hours, 5 hours, and 2 hours, respectively to complete the 3D generation. In contrast, DreamFusion only necessitates 1.5 hours since it skips the additional computations on the diffusion model.
@@
Another line of research uses the original Stable Diffusion to approximate the rendered image distribution. CSD [YGL23] and NFSD [KPCOL23] obtain an empirical solution by using negative prompt conditioned posterior of Stable Diffusion to represent the rendered image distribution. To accommodate the dynamic characteristics of generation, they use iteration steps related weights $\lambda_{s}$ and timesteps related weights $\lambda_{t}$ separately. ISM [LYL23] uses DDIM inversion [SME20] to get the score of rendered images. SSD [TWWZ23] applies a closed-form solution $r(\mathbf{x}_{t},\epsilon,t,y)$ to project 3D reprentations to the generated image domain of diffusion models. In Consistent3D [WZY24], the score of Stable Diffusion in smaller timestep $t_{2}$ is treated as the rendered image distribution score. Although this line of research approximates rendered image distribution in a training-free way, the final results may not always be satisfying. Since rendered images during the optimization process are commonly changeable and unnatural, being out-of-domain for the generated image distribution modeled by diffusion models, taking a precise negative prompt or applying a closed-form modeling strategy is unable to model the rendered image generation correctly. Worse still, an additional inference at each optimization iteration is introduced due to negative prompts (e.g., CSD [YGL23] and NFSD [KPCOL23]).

### v3

@@ Abstract
Score Distillation Sampling (SDS) has emerged as a prevalent technique for text-to-3D generation, enabling 3D content creation by distilling view-dependent information from text-to-2D guidance. However, they frequently exhibit shortcomings such as over-saturated color and excess smoothness. In this paper, we conduct a thorough analysis of SDS and refine its formulation, finding that the core design is to model the distribution of rendered images. Following this insight, we introduce a novel strategy called Variational Distribution Mapping (VDM), which expedites the distribution modeling process by regarding the rendered images as instances of degradation from diffusion-based generation. This special design enables the efficient training of variational distribution by skipping the calculations of the Jacobians in the diffusion U-Net. We also introduce timestep-dependent Distribution Coefficient Annealing (DCA) to further improve distilling precision. Leveraging VDM and DCA, we use Gaussian Splatting as the 3D representation and build a text-to-3D generation framework. Extensive experiments and evaluations demonstrate the capability of VDM and DCA to generate high-fidelity and realistic assets with optimization efficiency. {CCSXML} <ccs2012> <concept> <concept_id>10010147.10010371.10010382</concept_id> <concept_desc>Computing methodologies Image manipulation</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010396</concept_id> <concept_desc>Computing methodologies Shape modeling</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
@@ 1 Introduction
There has been an increasing need for digital 3D assets in video games [ZW12], mixed reality storytelling [CT14, HWW21], and digital fabrications [Ger12]. However, this widespread growth highlights the need for more efficient 3D content creation, as crafting 3D assets in graphics software is often time-consuming and labor-intensive. Thanks to the rapid development of diffusion models [HJA20, DN24], recent text-to-3D techniques offer a more accessible and efficient solution for controllable 3D content creation.
Based on the pre-trained text-to-image models [RBL22, SME21], Score Distillation Sampling (SDS) [PJBM23, WDL23] provides pixel-level guidance of 3D assets optimization and has become quite effective and popular for text-to-3D generation. Follow-up works further refine the supervision loss and optimization procedures to produce more realistic and semantic-aligned 3D models [WLW23, CCH23, CCJJ23, YGL23, KPCOL24, WZY24, LSS24, AKS24, WZSZ23, LYL23, TWWZ23, YCC23].
In this paper, we review SDS-based 3D generation and its variations, leading to a unifying goal across previous work on refining SDS formulation: establishing a variational distribution for rendered images. Our review and analysis demonstrate that prior efforts have predominantly adopted one of two strategies: 1) utilizing unmodified diffusion models to approximate the variational distribution [YGL23, KPCOL24, LYL23, WZY24] with improved loss designs, and 2) constructing the variational distribution directly through fine-tuning diffusion models [WLW23, LSS24, WZSZ23, YCC23] or pre-training a bespoke model [AKS24]. However, the former approach struggles to represent the distribution accurately, since the rendered images during the optimization process are out-of-domain (OOD) cases as to diffusion models [KPCOL24]. Meanwhile, the latter choice demands substantial resources and may exhibit instability during optimization.
To address these issues, we propose variational distribution mapping (VDM), a novel method treating rendered images as a degraded form of images generated by the diffusion model. VDM efficiently formulates the variational distribution of rendered images by modeling and optimizing the degradation process (i.e., lightweight trainable neural network), which enables us to establish a distribution mapping between the generated and the rendered image distributions (Figure 1). Compared with previous methods, VDM has two notable advantages: 1) By introducing the trainable degradation process, it eliminates the need for the complex Jacobian matrix calculations in the UNet of diffusion model, unlike previous variational distribution modeling approaches [WZSZ23, YCC23]; 2) VDM dynamically models the variational distribution of rendered images, surpassing methods that rely on an unmodified diffusion model with extra batch processing [KPCOL24, YGL23] for distribution estimation, thereby enhancing generation quality.
@@ 2.1 Text-to-3D Generation
With the rapid advancement of diffusion models and 3D representations, significant progress has been made in text-based generation. DreamFields [JMB22] introduced the optimization of Neural Radiance Fields (NeRF) using guidance from pre-trained CLIP models, but the quality was limited due to the CLIP loss’s insufficient semantic guidance. Building on this, DreamFusion [PJBM23] and SJC [WDL23] developed SDS, a novel loss function based on probability density distillation that provides pixel-level guidance by seeking specific modes in a text-guide diffusion model, enhancing the quality and efficiency of optimization-based 3D generation.
The advent of SDS marked a turning point that inspired a series of subsequent works in 3D generation [WLW23, YGL23, KPCOL24, LSS24, LYL23, YCC23, WZSZ23]. These studies have refined the generation of 3D assets from various perspectives. Some utilizes advanced differentiable 3D representations to enhance outcomes [LGT23, CCJJ23], while others [WZY24]focus on mitigating the “Janus” problem [HAK23, AKS24]. Zero-1-to-3 [LWH23] and MVDream [SWY24] fine-tuned the diffusion model with multi-view image datasets to improve 3D consistency. Perp-Neg [ASZ23] alleviates the Janus problem with view-dependent negative prompts, and ESD [WXF23] introduces a view-conditioned loss to improve multi-view generation.
A central emphasis in recent research lies in refining the original SDS methodology, aiming to secure more precise guidance from diffusion models. ProlificDreamer [WLW23] introduces a Variational Score Distillation Sampling technique to alleviate the mode collapse issue inherent to SDS. To further mitigate over-saturation and over-smoothness problems, TextMesh [TMT24] enforces textured optimizations with high SDS gradients only and integrates multi-view consistent diffusion. Make-it-3D [TWZ23] proposes two-stage optimizations to enhance textured appearance; Fantasia3D [CCJJ23] dynamically adjusts the time-dependent weighting function in SDS computations. Meanwhile, CSD [YGL23] and NFSD [KPCOL24] demonstrate the pivotal role of the conditional term in SDS for 3D generation. They employ additional negative prompts to refine the optimization process. Their explorations highly inspired our work, and we propose further improving SDS with the trainable construction of variational distribution. We will continue more in-depth discussions in Sec. 4.1.
@@ 2.2 Differentiable 3D Representations
The differentiable 3D representations are pivotal to text-to-3D generation. A 3D model parameterized by $\theta$ is rendered into an image from camera viewpoint $c$ using the differentiable rendering function $g(\theta,c)$. This allows for optimizing the 3D models through backpropagation to align with pixel-level guidance derived from diffusion models. Numerous differentiable 3D representations have been utilized in prior text-to-3D works [PJBM23, LGT23, MRP22, GAG23, TRZ23], including NeRF [MST20] which is frequently employed in generation tasks. However, NeRF’s volume rendering is computationally intensive, posing challenges for rendering high-resolution images efficiently enough to leverage the diffusion model’s guidance.
To mitigate this constraint, DMTet [SGY21] has been introduced to integrate both explicit and implicit representations, which has gained increasing utilizations [CCJJ23, LWW23]. Additionally, there is a growing interest in applying purely explicit representations to facilitate smooth shape manipulations within graphics software. This involves a gradient-based mesh deformation approach to create 3D models [GAG23, WMC24]. Notably, 3D Gaussian Splatting [KKLD23] has emerged as an efficient and high-quality explicit representation for reconstruction tasks and has been incorporated into several text-to-3D generation works [TRZ23, YFW24, LYL23]. We investigate Gaussian Splatting as the 3D representation for our generation framework.
3 Preliminaries
@@ 3.1 Diffusion Models
The Diffusion Model [HJA20, SME21, LYB24] is a likelihood-based generative model designed to approximate the data distribution $p_{data}$ starting from gaussian noise. Given a sample $\mathbf{x}$ from $p_{data}$, the model undergoes a forward diffusion process over a series of timesteps $t\in[0,t_{max}]$, whereby noise is incrementally added to transform $\mathbf{x}$ into Gaussian noise $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$. The diffusion at each step is described by:
@@
The training goal is to optimize $\mu_{\phi}(\mathbf{x}_{t})$ and $\Sigma_{\phi}(\mathbf{x}_{t})$ to maximize the log likelihood’s variational lower bound. In practice, the learning target is re-parameterized to the added noise $\epsilon\sim\mathcal{N}(0,I)$ which is used to produce $\mathbf{x}_{t}$ from $\mathbf{x}=\mathbf{x}_{0}$, and the notation of UNet in diffusion models is $\epsilon_{\phi}(\mathbf{x}_{t},t)$. The learning process of diffusion models can be interpreted as predicting the noise $\epsilon$ that corrupts the data $\mathbf{x}$. Furthermore, recent studies [LYB24, SE19] have shown that $\epsilon_{\phi}(\mathbf{x}_{t},t)$ corresponds to the score function $\nabla_{\mathbf{x}_{t}}\log p_{\phi}(\mathbf{x}_{t})$, indicating the model’s ability to guide $\mathbf{x}_{t}$ towards regions of higher density within $p_{\phi}(\mathbf{x_{t}})$, which can be approximated as:
@@
Classifier Free Guidance: Compared with unconditional image generation, text-guided image generation [HS21, SCS22, RBL22] has a higher demand. Conditioned on the text prompt $y$, diffusion models accept it as another input for the diffusion process, denoted as $\epsilon_{\phi}(\mathbf{x}_{t},t,y)$, with related score function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t}|y)$. Classifier Free Guidance (CFG) [HS21] is utilized as an implicit classifier to get the textual guidance for image generation. It has a changeable hyperparameter named CFG scale, hereafter denoted as $s$, and the original prediction is changed as:
@@ 3.2 SDS with Differentiable Rendering
As demonstrated in Sec. 2.1, SDS [PJBM23, WDL23] is an optimization-driven 3D generative technique that leverages pre-trained 2D diffusion models. Let $\mathbf{x}_{0}=g(\theta,c)$ represent the rendered image derived from differentiable 3D representations, parameterized by $\theta$ and camera pose $c$. The distribution of noisy rendered images is then formulated as follows:
@@
SDS adopts text-conditioned noisy real image distribution $p_{\phi}(\mathbf{x}_{t}|y)$ represented by pre-trained Stable Diffusion [RBL22] and optimizes the parameter $\theta$ by minimizing the following KL divergence for all timestep $t$:
@@
where $\epsilon\sim\mathcal{N}(0,I)$ and the UNet Jacobian term $\frac{\partial\epsilon_{\phi}(\mathbf{x}_{t},t,y)}{\partial\mathbf{x}_{t}}$ is ignored based on the analysis in DreamFusion [PJBM23].
@@
SDS [PJBM23] $\epsilon_{\phi}(\mathbf{x}_{t},t,y)-\epsilon$ Distillation baseline
@@
NFSD [KPCOL24] $\epsilon_{\phi}(\mathbf{x}_{t},t,y)-\lambda_{t}\epsilon_{\phi}(\mathbf{x}_{t},t,y_{neg})$
@@ 4.1 Systematic Review of Advancements in SDS
which delineates two components: the unconditional and the conditional terms. Numerous studies have indicated that the conditional term pivotal in ensuring the rendered image aligns with textual semantics [WLW23, YGL23, KPCOL24, TWWZ23]. This term is further guided by the score function $\nabla_{\mathbf{x}_{t}}\log p_{\phi}(y|\mathbf{x}_{t})$, as indicated in Eq. (5). Consequently, subsequent research mostly focuses on refining the unconditional term to improve SDS.
@@
Numerous studies have proposed various strategies to address the challenge of mode-seeking in 3D generation. Our review of recent advancements in refining the standard SDS approach reveals that these efforts generally focus on creating a variational distribution for the rendered images, which shifts the mode-seeking operation to minimizing the KL divergence between two distinct distributions. As illustrated in Table 1, VSD [WLW23] uses text-conditioned particles (or rendered images when the number of particles equals one) to build a variational distribution of 3D representations. VSD employs a LoRA [HysW22] model to swiftly optimize the score of rendered images, termed as $\epsilon_{\text{lora}}(\mathbf{x}_{t},t,y)$. DreamFlow [LSS24] follows VSD with a similar design, guided by theories of Schrodinger Bridge [LVH23]. ASD [WZSZ23] enhances the optimization methods of VSD within the framework of GAN [GPAM14], while LODS [YCC23] diversifies the modeling methods through learnable text embeddings denoted as $y_{\psi}$. LMC-SDS [AKS24] pre-trains a neural network to learn the score manifold corrective transitioning rendered images to generated images. The aforementioned studies collectively endeavor to determine the variational distribution of rendered images directly. Despite employing parameter-efficient modeling techniques, i.e., LoRA, the optimization of variational distribution requires additional computation of the UNet Jacobian for Stable Diffusion, which is time-consuming and prone to instability. For instance, given a text prompt, VSD, ASD, and LODs require approximately 8 hours, 5 hours, and 2 hours, respectively to complete the 3D generation. In contrast, DreamFusion only necessitates 1.5 hours since it skips the additional computations on the diffusion model.
@@
Another line of research uses the original Stable Diffusion to approximate the rendered image distribution. CSD [YGL23] and NFSD [KPCOL24] obtain an empirical solution by using negative prompt conditioned posterior of Stable Diffusion to represent the rendered image distribution. To accommodate the dynamic characteristics of generation, they use iteration steps related weights $\lambda_{s}$ and timesteps related weights $\lambda_{t}$ separately. ISM [LYL23] uses DDIM inversion [SME21] to get the score of rendered images. SSD [TWWZ23] applies a closed-form solution $r(\mathbf{x}_{t},\epsilon,t,y)$ to project 3D reprentations to the generated image domain of diffusion models. In Consistent3D [WZY24], the score of Stable Diffusion in smaller timestep $t_{2}$ is treated as the rendered image distribution score. Although this line of research approximates rendered image distribution in a training-free way, the final results may not always be satisfying. Since rendered images during the optimization process are commonly changeable and unnatural, being out-of-domain for the generated image distribution modeled by diffusion models, taking a precise negative prompt or applying a closed-form modeling strategy is unable to model the rendered image generation correctly. Worse still, an additional inference at each optimization iteration is introduced due to negative prompts (e.g., CSD [YGL23] and NFSD [KPCOL24]).